aline passing through (7,17,)and(5,y)is parrell to a line with the slope 6. What is tje value of y
Answer by checkley71(8403)

SLOPE=6
(Y2-Y1)/(X1-X1)=6
(Y-17)/(5-7)=6
(Y-17)/-2=6 NOW CROSS MULTIPLY
Y-17=-12
Y=-12+17
Y=5 ANSWER.
PROOF
(5-17)/(5-7)=6
-12/-2=6
6=6